### v1

# Causal Attention with Lookahead Keys

Abstract
In standard causal attention, each token’s query, key, and value (QKV) are static and encode only preceding context. We introduce CAuSal aTtention with Lookahead kEys (CASTLE), an attention mechanism that continually updates each token’s keys as the context unfolds. We term these updated keys lookahead keys because they belong to earlier positions yet integrate information from tokens that appear later relative to those positions, while strictly preserving the autoregressive property. Although the mechanism appears sequential, we derive a mathematical equivalence that avoids explicitly materializing lookahead keys at each position and enables efficient parallel training. On language modeling benchmarks, CASTLE consistently outperforms standard causal attention across model scales, reducing validation perplexity and improving performance on a range of downstream tasks.
[…]
Quanquan Gu at
1 Introduction
Causal attention [25] is a cornerstone of autoregressive sequence modeling, allowing each token to condition on its past while preserving the autoregressive structure that underpins language generation. Building on this mechanism, large language models (LLMs) have transformed natural language processing by scaling up the model size and the number of tokens trained [18, 3, 11]. While this trend has delivered impressive capabilities, it increasingly runs up against a practical bottleneck, that is high-quality tokens. This reality makes it imperative to improve the attention layer itself and to develop model architectures that are more token-efficient, delivering stronger performance under fixed training-token budgets.
In standard causal attention, each token’s query, key, and value (QKV) are computed from that token’s representation and remain fixed; they cannot incorporate information from subsequent tokens. As a result, a token’s QKV can encode only its preceding context. Recent work shows that such causal mask, which blocks each token’s access to its future information, limits models’ ability to capture global context, and impairs natural language understanding [13, 8, 21, 26, 30]. Here are some vivid illustrations that give intuitions for the limitations of causal masking. Garden-path sentences [17] are structurally ambiguous, often inducing an incorrect initial parse. For example, “the old man the boat”. Because garden-path sentences’ correct interpretation typically depends on information that appears later in the sentence, the causal mask restricting tokens to past context can cause models to struggle in resolving such ambiguities effectively [1]. In many tasks, the question/focus appears at the end of the input. Without access to future context, earlier tokens cannot encode the relevant information needed to anticipate the question/focus. As a result, early token representations may fail to capture important cues and global context dependencies [26].
Several recent works have proposed remedies to address these limitations on sentence embedding. BeLLM (Backward Dependency Enhanced LLM) [13] introduces backward dependencies which enables sentence embeddings to incorporate both past and future context, significantly improving performance on semantic textual similarity tasks and downstream applications. Echo Embeddings [21] duplicate the input and extracting representations from the second occurrence. The model effectively allows early tokens to attend to later ones and thus the model improves over classical LM embeddings without changing the architecture or requiring fine-tuning. Re-Reading (RE2) [26] enhances reasoning by prompting decoder-only LLMs to process the same question twice. The second pass provides tokens with access to global information from the first pass, capturing global dependencies and yielding stronger performance on arithmetic, commonsense, and symbolic reasoning tasks. Despite these advances, it remains unclear whether the above methods can benefit pretraining.
To address the shortcomings of standard causal attention in pretraining, we propose a novel attention mechanism, CAuSal aTtention with Lookahead kEys (CASTLE). In this approach, when generating the $(t+1)$-th token, we update keys of all preceding tokens so that keys of token $s$ ($1\leq s\leq t$) are able to additionally encode information from token $s+1$ to $t$. These keys are called lookahead keys. This design preserves the autoregressive structure while allowing the keys to evolve as the context unfolds. In Figure 1, we give an illustration of receptive fields of the keys in CASTLE. Although the mechanism appears recurrent, we establish a mathematical equivalence that avoids explicitly materializing the lookahead keys and enables efficient parallel training. We evaluate our approach on language modeling across multiple model scales. Experimental results show that CASTLE consistently outperforms standard causal attention in terms of validation perplexity and downstream task performance.
Figure 1: Receptive fields of keys in standard causal attention and CASTLE. The top row shows standard causal attention when generating the 4th token (left) and the 6th token (right). The bottom row shows CASTLE under the same settings. Here, key1, key2, $\cdots$ denote the keys corresponding to tokens 1, 2, $\cdots$, while $T_{1}$, $T_{2}$, $\cdots$ denote the tokens. In standard causal attention, keys are static: when generating the $(t+1)$-th token, each key $i$ with $1\leq i\leq t$ can only access information from $\left\{T_{1},\cdots,T_{i}\right\}$, and key $i$ remains the same for all later steps. In contrast, CASTLE continuously updates keys at each prediction step, i.e., when generating the $(t+1)$-th token, the receptive field of any key $i$ with $1\leq i\leq t$, is expanded to contain information from $\left\{T_{1},\cdots,T_{t}\right\}$.
The remainder of this paper is organized as follows. We discuss related work in Section 1.1. In Section 2.1, we elaborate on our motivations. In Section 2.2, we formally define CASTLE in its recurrent form. Section 2.3 proves an equivalent parallel formulation of CASTLE and develops efficient parallel training algorithms Section 2.4 introduces efficient inference algorithms together with the counterpart of the KV cache in CASTLE. Finally, Section 3 presents empirical results demonstrating the effectiveness of CASTLE across diverse tasks and model scales.
1.1 Related Work
Several studies have observed that the causal mask, by preventing tokens from accessing future information, can hinder a model’s ability to capture global context and thereby degrade its natural language understanding [13, 8, 21, 26, 30].
Much effort has been made to overcome this shortcoming in sentence embedding. BeLLM (Backward Dependency Enhanced LLM) [13] modifies decoder layers to enable sentence embeddings to integrate both past and future information. This yields substantial improvements on semantic textual similarity and downstream tasks. Echo Embeddings [21] duplicate the input sequence and extract embeddings from the second occurrence, letting early tokens attend to later ones without modifying the architecture. Similarly, Re-Reading (RE2) [26] prompts models to process inputs twice, so the second pass captures global information obtained in the first. These methods improve embedding quality and reasoning, but their benefits in large-scale pretraining remain unclear.
Selective Attention [12] introduces a parameter-free modification where tokens can mask out irrelevant or outdated tokens from future attention. By subtracting accumulated selection scores from the attention logits, selective attention reduces reliance on unneeded context. As a result, it achieves substantial memory and compute savings without degrading perplexity. However, selective attention primarily emphasizes filtering unneeded past tokens to enhance efficiency. As discussed in the introduction, many scenarios, such as garden-path sentences or cases where the key information appears at the end of the input, require mechanisms that actively incorporate crucial future information. Whether selective attention can address this challenge remains uncertain.
2 Causal Attention with Lookahead Keys
Our motivations are discussed in Section 2.1. Formal mathematical definitions of CASTLE in recurrent form are provided in Section 2.2. Direct application of the recurrent form of CASTLE in Section 2.2 is impractical for large-scale pretraining. To address this, we present efficient pretraining algorithms in Section 2.3. In Section 2.4, we describe efficient inference algorithms along with the counterparts of the KV cache in CASTLE.
2.1 Motivations
We first recall the standard causal attention. Given contextualized representations $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$ where $L$ is the sequence length and $d_{\rm model}$ is the hidden dimension. The standard causal attention first computes $\boldsymbol{\mathit{Q}}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}_{Q}$, $\boldsymbol{\mathit{K}}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}_{K}$, $\boldsymbol{\mathit{V}}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}_{V}\in\mathbb{R}^{L\times d}$, where $d$ is the head dimension. Then, the standard causal attention is computed as follows
$\begin{split}\text{CausalAttention}(\boldsymbol{\mathit{X}}^{L})=\text{row\_softmax}\Big{(}\frac{\boldsymbol{\mathit{Q}}\boldsymbol{\mathit{K}}^{\top}}{\sqrt{d}}+\boldsymbol{\mathit{M}}^{C}\Big{)}\boldsymbol{\mathit{V}}\in\mathbb{R}^{L\times d},\end{split}$ (1)
where $\boldsymbol{\mathit{M}}^{C}\in\mathbb{R}^{L\times L}$ is the causal mask which prevents each token from attending to its future tokens, i.e., $\boldsymbol{\mathit{M}}^{C}_{ij}=0$ if $i\geq j$ and $\boldsymbol{\mathit{M}}^{C}_{ij}=$-$\infty$ otherwise.
To explain our motivations, we begin with the recurrent form of standard causal attention. Consider the case when generating the $(t+1)$-th token. Given contextualized representations $\boldsymbol{\mathit{X}}^{t}=\begin{pmatrix}\boldsymbol{\mathit{x}}_{1}\\ \boldsymbol{\mathit{x}}_{2}\\ \vdots\\ \boldsymbol{\mathit{x}}_{t}\end{pmatrix}\in\mathbb{R}^{t\times d_{\rm model}}$, where $\boldsymbol{\mathit{x}}_{s}$ is the representation of token $s$. Unless otherwise specified, all vectors in this paper are treated as row vectors rather than column vectors.
The query, key and value of token $s$ are $\boldsymbol{\mathit{q}}_{s}=\boldsymbol{\mathit{x}}_{s}\boldsymbol{\mathit{W}}_{Q}$, $\boldsymbol{\mathit{k}}_{s}=\boldsymbol{\mathit{x}}_{s}\boldsymbol{\mathit{W}}_{K}$, $\boldsymbol{\mathit{v}}_{s}=\boldsymbol{\mathit{x}}_{s}\boldsymbol{\mathit{W}}_{V}$. We also denote $\boldsymbol{\mathit{K}}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{K}$ and $\boldsymbol{\mathit{V}}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{V}$. Then, the standard causal attention follows the recurrent form
$\begin{split}\text{causal-attention}(\boldsymbol{\mathit{X}}^{t})={\rm softmax}\left(\frac{\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{K}}_{t}}^{\top}}{\sqrt{d}}\right)\boldsymbol{\mathit{V}}_{t}=\frac{\sum_{s=1}^{t}\exp\big{(}\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{k}}_{s}}^{\top}/\sqrt{d}\big{)}\boldsymbol{\mathit{v}}_{s}}{\sum_{s=1}^{t}\exp\big{(}\boldsymbol{\mathit{q}}_{t}{\boldsymbol{\mathit{k}}_{s}}^{\top}/\sqrt{d}\big{)}}\in\mathbb{R}^{1\times d}.\end{split}$
Due to the autoregressive structure, each $\boldsymbol{\mathit{x}}_{s}$ only encodes information from token $1$ to $s$. Thus, when generating token $t+1$ with $t+1>s$, each $\boldsymbol{\mathit{k}}_{s}$ only contains information from token $1$ to $s$ without containing information from token $s+1$ to $t$. This can impair models’ ability of natural language understanding, yielding high-quality text embedding and capturing global context as mentioned in the introduction.
This motivates us to propose a novel attention mechanism, causal attention with lookahead keys (CASTLE), i.e., when generating the $(t+1)$-th token, we first update keys $\boldsymbol{\mathit{k}}_{s}$ of preceding tokens $s$ with $s<t+1$ to additionally incorporate information from token $s+1$ to $t$. We refer to these as lookahead keys because their representations renew with the growing context. In this way, lookahead keys may lead to more accurate attention scores while preserving the autoregressive property of the model.
Before describing the details of this mechanism, we first answer the following questions.
$\bullet$ Why do we use lookahead keys instead of lookahead queries? The answer parallels the reason why key–value (KV) pairs are cached instead of queries (Q). Each $\boldsymbol{\mathit{q}}_{s}$ is only used once, namely when generating token $s+1$. Because we are designing an autoregressive model, past queries cannot be modified after generation, making it meaningless to update $\boldsymbol{\mathit{q}}_{s}$. In contrast, $\boldsymbol{\mathit{k}}_{s}$ is multiplied by the queries $\boldsymbol{\mathit{q}}_{t}$ of all subsequent tokens $t\geq s$. Keeping $\boldsymbol{\mathit{k}}_{s}$ updated therefore can benefit all later tokens by possibly producing more accurate attention scores.
We also remark that renewing the values $\boldsymbol{\mathit{v}}_{s}$ in a similar way to what we do for lookahead keys could potentially be beneficial. In this work, however, we focus on updating keys because doing so admits a mathematical equivalence (shown in Section 2.3) with $O(L^{2}d)$ computational complexity. Designing efficient parallel training algorithms that utilize lookahead values remains an interesting direction for future research.
$\bullet$ How do we maintain the model autoregressive with lookahead keys? When generating token $t+1$, we update keys of each preceding token $s<t+1$ with information from token $s+1$ to $t$. Thus, all keys only contain information from token $1$ to $t$. Queries and values are naturally only containing information from tokens up to $t$. No future information from tokens $k>t$ is used. Thus, the model maintains autoregressive property.
Further details of our design are presented in Section 2.2.
2.2 Mathematical Definition in Recurrent Form
We give mathematical definitions of CASTLE in this section. Let $L$ denote sequence length and $_{\rm model}, dd$ denote the hidden dimension and head dimension, respectively.
Throughout Section 2.2, we fix $t\in\left\{1,2,\cdots,L\right\}$ and consider the setting where $t$ tokens have already been generated and the model is about to generate the $(t+1)$-th token.
Denote the input contextualized representations $\boldsymbol{\mathit{X}}^{t}=\begin{pmatrix}\boldsymbol{\mathit{x}}_{1}\\ \boldsymbol{\mathit{x}}_{2}\\ \vdots\\ \boldsymbol{\mathit{x}}_{t}\end{pmatrix}\in\mathbb{R}^{t\times d_{\rm model}}$, where $\boldsymbol{\mathit{x}}_{s}$ is the representation of token $s$.
Utilizing lookahead keys lies at the core of CASTLE. However, the way a model learns to encode information into keys of token $s$ from past tokens ($k\leq s$) may differ from how it encodes information from subsequent tokens (tokens $s<k<t+1$) when generating the $(t+1)$-th token. To address this, we adopt a hybrid design. Specifically, we keep half of the keys the same as in standard causal which we call causal keys, while allowing the remaining half to renew as the context progresses which we call lookahead keys.
For each preceding token $s$ ($1\leq s<t+1$), causal keys of token $s$ is a projection of $\boldsymbol{\mathit{x}}_{s}$, while lookaheads keys of token $s$ contain information from representations $\left\{\boldsymbol{\mathit{x}}_{s+1},\cdots,\boldsymbol{\mathit{x}}_{t}\right\}$. The receptive fields of causal keys and lookahead keys are illustrated in Figure 2.
We first project $\boldsymbol{\mathit{X}}^{t}$ into key and value matrices $\boldsymbol{\mathit{K}}^{U}_{t}$, $\boldsymbol{\mathit{V}}^{U}_{t}$, $\boldsymbol{\mathit{K}}^{C}_{t}$, $\boldsymbol{\mathit{V}}^{C}_{t}\in\mathbb{R}^{t\times d}$ by
$\begin{split}\boldsymbol{\mathit{K}}^{U}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{K}^{U},\boldsymbol{\mathit{V}}^{U}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{V}^{U},\end{split}$
and
$\begin{split}\boldsymbol{\mathit{K}}^{C}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{K}^{C},\boldsymbol{\mathit{V}}^{C}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}_{V}^{C}\end{split}$
as well as query matrix $\boldsymbol{\mathit{Q}}^{U}_{t}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}^{U}_{Q}\in\mathbb{R}^{t\times d}$ and query vector $\boldsymbol{\mathit{q}}^{C}_{t}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{C}_{Q}\in\mathbb{R}^{1\times d}$. Here, $\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{W}}^{U}_{K}$, $\boldsymbol{\mathit{W}}^{U}_{V}$, $\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{W}}^{C}_{K}$, $\boldsymbol{\mathit{W}}^{C}_{V}\in\mathbb{R}^{d_{\rm model}\times d}$ are learnable matrices.
The matrices $\boldsymbol{\mathit{K}}^{U}_{t}$, $\boldsymbol{\mathit{V}}^{U}_{t}$, $\boldsymbol{\mathit{Q}}^{U}_{t}$ are used to generate the lookahead key $\boldsymbol{\mathit{U}}^{t}$. Then, the causal key $\boldsymbol{\mathit{K}}^{C}_{t}$ and the lookahead key $\boldsymbol{\mathit{U}}^{t}$ are multiplied by the query vector $\boldsymbol{\mathit{q}}^{C}_{t}$ to get the attention scores. Then, $\boldsymbol{\mathit{V}}^{C}_{t}$ are multiplied by the attention weights to get the output. Before we elaborate on details in the definition of CASTLE, we first give formal definitions of causal keys and lookahead keys.
Figure 2: Receptive fields of causal keys and lookahead keys with respect to contextualized representations and tokens (excluding the first layer) when generating the 6th token. Tokens are denoted by $T_{i}$ and their contextualized representations by $\boldsymbol{\mathit{x}}_{i}$. Causal key $i$ corresponds to the causal key of token $i$, while lookahead key $i$ corresponds to the lookahead key of token $i$. When generating the $(t+1)$-th token, for token $s$ ($s<t+1$), the causal key of token $s$ is a projection of $\boldsymbol{\mathit{x}}_{s}$. Due to the softmax in attention, except in the first layer, causal keys of $s$ attend over tokens $\left\{T_{1},\cdots,T_{s}\right\}$. For token $s<t$, lookahead keys of $s$ incorporate information from $\left\{\boldsymbol{\mathit{x}}_{s+1},\cdots,\boldsymbol{\mathit{x}}_{t}\right\}$ and attend over all existing tokens $\left\{T_{1},\cdots,T_{t}\right\}$. Since the last row of $\boldsymbol{\mathit{M}}^{U}$ is defined as $\boldsymbol{\mathit{M}}^{U}_{t}]_{t,:}=(-\infty)^{1\times t}$, lookahead keys of token $t$ are all-zeros vectors and thus have empty receptive fields when generating the $(t+1)$-th token.
Causal Keys. The causal keys in CASTLE is defined similarly to the keys in standard causal attention. More specifically, causal keys are defined as
$\begin{split}\boldsymbol{\mathit{K}}^{C}_{t}=\begin{pmatrix}\boldsymbol{\mathit{k}}^{C}_{1}\\ \boldsymbol{\mathit{k}}^{C}_{2}\\ \vdots\\ \boldsymbol{\mathit{k}}^{C}_{t}\end{pmatrix}=\boldsymbol{\mathit{X}}^{t}\boldsymbol{\mathit{W}}^{C}_{K}\in\mathbb{R}^{t\times d}.\end{split}$
The $s$-th row $\boldsymbol{\mathit{k}}_{s}$ of $\boldsymbol{\mathit{K}}^{C}_{t}$ satisfying $\boldsymbol{\mathit{k}}_{s}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{C}_{K}$ is the causal key of token $s$. Causal keys are static, i.e., the $s$-th rows of $\boldsymbol{\mathit{K}}^{C}_{t}$ and $\boldsymbol{\mathit{K}}^{C}_{t^{\prime}}$ are equal to each other whenever $t,t^{\prime}\geq s$.
Lookahead Keys. We utilize a structure similar to the attention mechanism to define lookahead keys.
More specifically, the lookahead keys are defined as
$\begin{split}\boldsymbol{\mathit{U}}^{t}=\begin{pmatrix}\boldsymbol{\mathit{u}}^{t}_{1}\\ \boldsymbol{\mathit{u}}^{t}_{2}\\ \vdots\\ \boldsymbol{\mathit{u}}^{t}_{t}\end{pmatrix}={\rm sigmoid}\big{(}\frac{\boldsymbol{\mathit{Q}}^{U}_{t}{\boldsymbol{\mathit{K}}^{U}_{t}}^{\top}}{\sqrt{d}}+\boldsymbol{\mathit{M}}^{U}_{t}\big{)}\boldsymbol{\mathit{V}}^{U}_{t}\in\mathbb{R}^{t\times d},\end{split}$ (2)
where $\boldsymbol{\mathit{M}}^{U}_{t}\in\mathbb{R}^{t\times t}$ is a mask matrix with $\boldsymbol{\mathit{M}}^{U}_{t}]_{ij}=0$ if $i<j$ and $\boldsymbol{\mathit{M}}^{U}_{t}]_{ij}=-\infty$ otherwise. The $s$-th row $\boldsymbol{\mathit{u}}^{t}_{s}$ of $\boldsymbol{\mathit{U}}^{t}$ is the lookahead key of token $s$. We remark that in $\boldsymbol{\mathit{u}}^{t}_{s}$, the superscript $t$ indicates that $\boldsymbol{\mathit{u}}^{t}_{s}$ is defined when $t$ tokens have already been generated and we are about to generate the $(t+1)$-th token, while the subscript $s$ indicates $\boldsymbol{\mathit{u}}^{t}_{s}$ is the $s$-th row of $\boldsymbol{\mathit{U}}^{t}$.
The definition of $\boldsymbol{\mathit{M}}^{U}_{t}$ guarantees that the lookahead key of token $s$, $\boldsymbol{\mathit{u}}^{t}_{s}$, is exposed to information from $\left\{\boldsymbol{\mathit{x}}_{s+1},\cdots,\boldsymbol{\mathit{x}}_{t}\right\}$. Since $\boldsymbol{\mathit{u}}^{t}_{s}$ keeps renewing as the context goes, it is natural that $\boldsymbol{\mathit{u}}^{t}_{s}\neq\boldsymbol{\mathit{u}}^{t+1}_{s}$.
An illustration for lookahead keys can be found in Figure 3. We have the following remarks regarding the definition of lookahead keys in (2).
Figure 3: Illustration for the definition of lookahead keys in (2) when generating the $4$-th token. Let $d_{\rm model}$ and $d$ denote the hidden dimension and head dimension, respectively. In this figure, we set $t=3$ and $d=2$.
  • Why are we using sigmoid? The $\rm sigmoid$} function is used in (2) instead of softmax due to the consideration that when generating token $t+1$, for token $s$ with $s<t+1$, synthesizing information contained in tokens $s+1$ to $t$ should not be compulsory. However, since the probabilities in softmax sum up to $1$, which forces $\boldsymbol{\mathit{u}}^{t}_{s}$ to incorporate information from tokens $s+1$ to $t$ and is not desired.
  • Lookahead keys $\boldsymbol{\mathit{U}}^{t}$ maintains autoregressive property. First, CASTLE is defined in a recurrent form which is naturally autoregressive. Second, when generating the $(t+1)$-th token, each $\boldsymbol{\mathit{u}}^{t}_{s}$ is only exposed to information from representations of tokens $s+1$ to $t$ as in (2). No information from tokens which are not yet generated is exposed.
  • Lookahead keys $\boldsymbol{\mathit{U}}^{t}$ only occur in CASTLE’s recurrent form definition and inference, but cannot be materialized in parallel training. Since $\boldsymbol{\mathit{u}}^{t}_{s}$ and $\boldsymbol{\mathit{u}}^{t+1}_{s}$ may vary, this prevents us from materializing $\boldsymbol{\mathit{U}}^{t}$ for each $t$. The computation cost in (2) is $O(t^{2}d)$. If we materialize all $\boldsymbol{\mathit{U}}^{t}$ in parallel, the computational cost is at least $\sum_{t=1}^{L}t^{2}d=O(L^{3}d)$ which makes training on large-scale datasets impractical. In Section 2.3, we will give an equivalent form which removes the need of materializing each $\boldsymbol{\mathit{U}}^{t}$ and enables efficient parallel training.
Figure 4: Illustration of CASTLE in recurrent form when generating the 4th token. The causal and lookahead keys are queried by $\boldsymbol{\mathit{q}}^{C}_{t}$ to generate their respective attention scores, which are combined and then goes through softmax to yield attention weights $\boldsymbol{\mathit{p}}_{t}$. These weights are then multiplied by the value matrix $\boldsymbol{\mathit{V}}^{C}_{t}$ to compute the output $\boldsymbol{\mathit{o}}_{t}=\text{attention}\left(\boldsymbol{\mathit{X}}^{t}\right)\in\mathbb{R}^{1\times d}$
CASTLE in Recurrent Form. After defining causal keys and lookahead keys, we are ready to give the formula of CASTLE in recurrent form. To generate the $(t+1)$-th token, we utilize both the causal keys $\boldsymbol{\mathit{K}}^{C}_{t}\in\mathbb{R}^{t\times d}$ and the lookahead keys $\boldsymbol{\mathit{U}}^{t}$. More specifically, let the causal-key attention scores be
$\begin{split}\boldsymbol{\mathit{s}}^{C}_{t}=\frac{\boldsymbol{\mathit{q}}^{C}_{t}{\boldsymbol{\mathit{K}}^{C}_{t}}^{\top}}{\sqrt{d}}\in\mathbb{R}^{1\times t}.\end{split}$ (3)
Let the lookahead-key attention scores be
$\begin{split}\boldsymbol{\mathit{s}}^{U}_{t}={\frac{\boldsymbol{\mathit{q}}^{C}_{t}{\boldsymbol{\mathit{U}}^{t}}^{\top}}{\sqrt{d}}}\in\mathbb{R}^{1\times t}.\end{split}$ (4)
Then, we define attention weights by combining the above attention scores as follows
$\begin{split}\boldsymbol{\mathit{p}}_{t}={\rm softmax}\left(\boldsymbol{\mathit{s}}^{C}_{t}-{\rm SiLU}\left(\boldsymbol{\mathit{s}}^{U}_{t}\right)\right)\in\mathbb{R}^{1\times t},\end{split}$ (5)
where ${\rm SiLU}(\boldsymbol{\mathit{x}})=\boldsymbol{\mathit{x}}\odot{\rm sigmoid}(\boldsymbol{\mathit{x}})$.
Then, the output is calculated as
$\begin{split}\text{attention}\left(\boldsymbol{\mathit{X}}^{t}\right)=\boldsymbol{\mathit{p}}_{t}\boldsymbol{\mathit{V}}^{C}_{t}\in\mathbb{R}^{1\times d}.\end{split}$ (6)
An illustration of CASTLE in its recurrent form can be found in Figure 4.
We remark that $\rm SiLU$} is applied in (5) because our ablation study shows that $\boldsymbol{\mathit{p}}_{t}={\rm softmax}\left(\boldsymbol{\mathit{s}}^{C}_{t}-{\rm SiLU}\left(\boldsymbol{\mathit{s}}^{U}_{t}\right)\right)$ outperforms $\boldsymbol{\mathit{p}}_{t}={\rm softmax}\left(\boldsymbol{\mathit{s}}^{C}_{t}-\boldsymbol{\mathit{s}}^{U}_{t}\right)$ in downstream evaluation. We hypothesize that this improvement arises because many past tokens become “noise” as the context grows. The $\rm SiLU$} transformation effectively acts as a gate, regulating the degree to which each past token should be forgotten.
2.3 Efficient Parallel Training
In this section, we introduce our efficient parallel training algorithm. As discussed in Section 2.2, a straightforward materializing each $\boldsymbol{\mathit{U}}^{t}$ in parallel incurs at least $O(L^{3}d)$ computational costs. Such complexity makes training on large-scale datasets infeasible. To address this, we derive an equivalent parallel formulation of CASTLE that eliminates the need to materialize lookahead keys. This formulation then serves as the basis for designing an efficient algorithm.
CASTLE in Parallel Form. Let $\text{attention}\left(\boldsymbol{\mathit{X}}^{t}\right)\in\mathbb{R}^{1\times d}$ denote the output when generating the $(t+1)$-th token as in (6). Then, given the inputs $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$, the concatenated outputs are denoted by
$\begin{split}\text{Attention}(\boldsymbol{\mathit{X}}^{L})=\begin{pmatrix}\text{attention}(\boldsymbol{\mathit{X}}^{1})\\ \text{attention}(\boldsymbol{\mathit{X}}^{2})\\ \vdots\\ \text{attention}(\boldsymbol{\mathit{X}}^{L})\end{pmatrix}\in\mathbb{R}^{L\times d_{\rm model}}.\end{split}$ (7)
The following theorem provides a parallel formulation of CASTLE that is equivalent to the recurrent form introduced in Section 2.2, but eliminates the need to explicitly materialize lookahead keys.
Theorem 1.
[…]
Then, the outputs $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$ as in (7) satisfies that
[…]
Here, $\boldsymbol{\mathit{M}}^{C}$, $\widetilde{\boldsymbol{\mathit{M}}}^{C}$ are the causal masks which prevent tokens from attending to their future tokens, i.e., $\boldsymbol{\mathit{M}}^{C}_{ij}=0$ if $i\geq j$ and $\boldsymbol{\mathit{M}}^{C}_{ij}=$-$\infty$ otherwise; $\widetilde{\boldsymbol{\mathit{M}}}^{C}_{ij}=1$ if $i\geq j$ and $\widetilde{\boldsymbol{\mathit{M}}}^{C}_{ij}=0$ otherwise. $\boldsymbol{\mathit{M}}^{U}$ is the same as in (2), with $\boldsymbol{\mathit{M}}^{U}_{ij}=0$ if $i<j$ and $\boldsymbol{\mathit{M}}^{U}_{ij}=$-$\infty$ otherwise.
The proof of Theorem 1 is deferred to Appendix 5.
Efficient Parallel Training. Theorem 1 establishes the equivalence between the recurrent and parallel formulations of CASTLE. However, computing $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$ directly from Theorem 1 still requires $\Omega(L^{3})$ complexity, since (8) involves matrix multiplications between $L\times L$ matrices.
[…]
Theorem 2.
Given $\boldsymbol{\mathit{X}}^{L}$’s query, key and value matrices $\boldsymbol{\mathit{Q}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{K}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{K}$, $\boldsymbol{\mathit{V}}^{U}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{V}$, $\boldsymbol{\mathit{Q}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{K}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{K}$, $\boldsymbol{\mathit{V}}^{C}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{V}$, Algorithm 1 (forward pass) and Algorithm 3 (backward pass) enables efficient parallel training and can compute $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$ and the gradients with computational complexity $O(L^{2}d)$ and space complexity $O(Ld)$.
The proof of Theorem 2 is given in Appendix 7.
2.4 Efficient Inference with UQ-KV Cache
In this section, we introduce the inference algorithm. We first introduce the decoding algorithm. The decoding algorithm consists of the following 2 steps: updating step and combining step. Fix $t\in\left\{1,2,\dots,L\right\}$, and consider generating the $(t+1)$-th token.
Updating step. We generate lookahead keys $\boldsymbol{\mathit{U}}^{t}$ in the updating step. First, we compute $\boldsymbol{\mathit{q}}^{U}_{t}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{k}}^{U}_{t}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{U}_{K}$ and $\boldsymbol{\mathit{v}}^{U}_{t}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{U}_{V}$. Next, Rather than computing $\boldsymbol{\mathit{U}}^{t}$ directly from (2), which requires $O(t^{2}d)$ computation, we update $\boldsymbol{\mathit{U}}^{t}$ recursively
[…]
The proof of (10) is given in Appendix 8. Next, we discuss the caching strategy and FLOPs of the updating step.
  • Caching in updating step. First, it is obvious that we need to cache $\boldsymbol{\mathit{U}}^{t}$ so that we can implement the recursive formula (10). Second, we need to cache $\boldsymbol{\mathit{Q}}^{U}_{t}$ because $\boldsymbol{\mathit{q}}^{U}_{s}$ is used in any update from $\boldsymbol{\mathit{U}}^{t-1}$ to $\boldsymbol{\mathit{U}}^{t}$ with $s\leq t-1$. As $\boldsymbol{\mathit{k}}^{U}_{t}$ and $\boldsymbol{\mathit{v}}^{U}_{t}$ are only used in the update from $\boldsymbol{\mathit{U}}^{t-1}$ to $\boldsymbol{\mathit{U}}^{t}$ and never used again in update from $\boldsymbol{\mathit{U}}^{j-1}$ to $\boldsymbol{\mathit{U}}^{j}$ with $j\neq t$, we do not need to cache any other variables except $\boldsymbol{\mathit{U}}^{t}$ and $\boldsymbol{\mathit{Q}}^{U}_{t}$ for the updating step.
  • FLOPs in updating step. With cached $\boldsymbol{\mathit{U}}^{t-1}$ and $\boldsymbol{\mathit{Q}}^{U}_{t-1}$, the updating formula (10) needs only $O(td)$ FLOPs. And computing $\boldsymbol{\mathit{q}}^{U}_{t}$, $\boldsymbol{\mathit{k}}^{U}_{t}$ and $\boldsymbol{\mathit{v}}^{U}_{t}$ needs $O(dd_{\rm model})$ FLOPs, yielding total FLOPs of $O(dd_{\rm model}+td)$.
Combining step. In the combining step, we compute $\boldsymbol{\mathit{q}}^{C}_{t}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{k}}^{C}_{t}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{C}_{K}$ and $\boldsymbol{\mathit{v}}^{C}_{t}=\boldsymbol{\mathit{x}}_{t}\boldsymbol{\mathit{W}}^{C}_{V}$. Next, the attention outputs are then obtained by applying (3), (4), (5) and (6) with $O(td)$ FLOPs. At this stage, since we already cached $\boldsymbol{\mathit{U}}^{t}$ in the updating step, only $\boldsymbol{\mathit{K}}^{C}_{t}$ and $\boldsymbol{\mathit{V}}^{C}_{t}$ need to be stored.
UQ-KV cache. From the above analysis, the counterpart of the KV cache in CASTLE consists of $\boldsymbol{\mathit{U}}^{t}$, $\boldsymbol{\mathit{Q}}^{U}_{t}$, $\boldsymbol{\mathit{K}}^{C}_{t}$, and $\boldsymbol{\mathit{V}}^{C}_{t}$, which we collectively refer to as the UQ-KV cache. All other variables, including $\boldsymbol{\mathit{q}}^{C}_{t}$, $\boldsymbol{\mathit{k}}^{U}_{t}$, and $\boldsymbol{\mathit{v}}^{U}_{t}$, can be safely disposed of after use.
For the prefilling stage, let the prompt length be $L$ and inputs $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$. We first compute $\boldsymbol{\mathit{Q}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{Q}$, $\boldsymbol{\mathit{K}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{K}$, $\boldsymbol{\mathit{V}}^{U}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{U}_{V}$, $\boldsymbol{\mathit{Q}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{Q}$, $\boldsymbol{\mathit{K}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{K}$, $\boldsymbol{\mathit{V}}^{C}_{L}=\boldsymbol{\mathit{X}}^{L}\boldsymbol{\mathit{W}}^{C}_{V}$. Then, we apply the forward pass of the efficient parallel training algorithm (Algorithm 1) to get $\text{Attention}(\boldsymbol{\mathit{X}}^{L})$. For the UQ-KV cache, since we already have $\boldsymbol{\mathit{Q}}^{U}_{L}$, $\boldsymbol{\mathit{K}}^{C}_{L}$ and $\boldsymbol{\mathit{V}}^{C}_{L}$, we only need to obtain $\boldsymbol{\mathit{U}}^{L}$. This can be done similarly to FlashAttention-2 [5] due to the similarity between (2) and standard causal attention. The complete prefilling procedure is given in Algorithm 4.
The analysis above leads to the following theorem.
Theorem 3.
Given inputs $\boldsymbol{\mathit{X}}^{L}\in\mathbb{R}^{L\times d_{\rm model}}$, prefilling has computational complexity $O(Ldd_{\rm model}+L^{2}d)$ and space complexity $O(Ld)$. During the decoding stage, when generating the $t$-th token, the computational complexity is $O(td+dd_{\rm model})$ and the UQ-KV cache requires $O(td)$ memory.
2.5 Multi-Head Causal Attention with Lookahead Keys
As in standard causal attention, we also utilize multi-head mechanism. Let $n$ denote the number of heads. In each head $i$, when generating the $t$-th token, denote the output as in (6) by $\text{attention}_{i}(\boldsymbol{\mathit{X}}^{t})\in\mathbb{R}^{1\times d}$. Then, the output of multi-head causal attention with lookahead keys (multi-head CASTLE) is calculated as
$\begin{split}\text{multi-head-attention}(\boldsymbol{\mathit{X}}^{t})=\text{Concat}\left(\text{attention}_{1}(\boldsymbol{\mathit{X}}^{t}),\dots,\text{attention}_{n}(\boldsymbol{\mathit{X}}^{t})\right)\boldsymbol{\mathit{W}}^{O}\in\mathbb{R}^{1\times d},\end{split}$ (11)
where $\boldsymbol{\mathit{W}}^{O}\in\mathbb{R}^{nd\times d_{\rm model}}$ is a learnable matrix. The formula for forward pass in parallel training and more details of multi-head CASTLE are in Appendix 6.
3 Experiments
3.1 Experimental Setup
Our baseline follows the improved Transformer architecture with SwiGLU [20], Rotary Positional Embeddings [22], and RMSNorm [29] following LLaMA [24]. We train models on four scales from scratch: small (0.16B), medium (0.35B), large (0.75B) and XL (1.3B). The medium, large and XL baseline models mirror the configuration of GPT-3 [3]. For the small setting, we increase the number of heads and the hidden dimension relative to the original GPT-3 configuration to better align parameter counts between standard causal attention and CASTLE. To isolate the effect of the attention mechanism, we replace standard causal attention with dynamic attention with evolving keys while keeping all other components unchanged for a fair comparison. We use the AdamW optimizer [14] and follow the training recipe of [6]. All models are trained on FineWeb-Edu dataset [15] for 50B tokens. Further details of experimental setup can be found in Appendix 9.
3.2 Training & Validation Loss
We report training and validation loss and perplexity in Table 1. Training loss and validation loss curves are shown in Figure 5, Figure 7, Figure 8 and Figure 9. CASTLE consistently outperforms the baseline models in both training and validation loss across all model scales.
Specifically, after training 50B tokens, CASTLE outperforms baselines across all model scales and achieves validation losses that are 0.0059, 0.0245, 0.0356, and 0.0348 lower than the baseline for the small, medium, large, and XL models, respectively. The performance gains are particularly significant in the medium, large, and XL models.
We hypothesize that this improvement stems from lookahead keys in CASTLE. By incorporating lookahead keys, CASTLE facilitates better global context capture. However, smaller models may struggle to fully leverage this mechanism due to limited capacity to model complex global relationships. As a result, the improvement margin for the small model is less significant compared to larger models.
Furthermore, as shown in Table 4, CASTLE has the same or fewer parameters than their baseline counterparts. This further underscores CASTLE’s effectiveness in improving model performance.
Table 1: Training and validation loss and perplexity for models with standard causal attention and CASTLE. We use “S”, “M”, “L”, “XL” to denote model scales. Each model is trained on FineWeb-Edu for 50B tokens. Significant improvements in loss or perplexity are in bold.
Train Eval
Loss PPL Loss PPL
Baseline-S 2.795 16.364 2.798 16.411
CASTLE-S 2.789 16.259 2.792 16.315
Baseline-M 2.641 14.030 2.639 14.004
CASTLE-M $\bm{2.616}$ $\bm{13.684}$ $\bm{2.615}$ $\bm{13.665}$
Baseline-L 2.513 12.346 2.507 12.269
CASTLE-L $\bm{2.476}$ $\bm{11.897}$ $\bm{2.472}$ $\bm{11.840}$
Baseline-XL 2.430 11.360 2.426 11.309
CASTLE-XL $\bm{2.401}$ $\bm{11.031}$ $\bm{2.391}$ $\bm{10.922}$
Figure 5: Training and validation loss curves of XL models. Training loss curve is smoothened with a moving window of 2000 training steps. Validation loss is evaluated every 100 training steps on 40M tokens, and its curve is smoothened by a moving window of 20 evaluation intervals. Loss curves for the small, medium and large models can be found in Figure 7, Figure 8 and Figure 9 of Appendix 9.3. After 50B training tokens, CASTLE-XL achieves a $\bm{0.0294}$ lower training loss and a $\bm{0.0348}$ lower validation loss compared to Baseline-XL.
3.3 Downstream Evaluation
We evaluate CASTLE on a diverse set of downstream benchmarks, including ARC [27], BoolQ [4], HellaSwag [28], MMLU [10], OBQA [16], PIQA [2], Winograde [19] using lm-evaluation-harness [9]. Table 1 reports normalized accuracy for ARC-Challenge, HellaSwag, OBQA, and PIQA, and standard accuracy for the other benchmarks. Results are reported in Table 2 (0-shot) and Table 3 (5-shot). Across all model scales and evaluation settings, CASTLE consistently outperforms the baseline in average downstream accuracy. In the 0-shot setting, CASTLE demonstrates robust improvements on ARC-E, BoolQ, HellaSwag, and OBQA across all scales, highlighting enhanced reasoning and commonsense capabilities. In the 5-shot setting, CASTLE further strengthens performance, with consistent gains on ARC-C across all scales, underscoring its strong few-shot generalization on reasoning and knowledge integration abilities. These findings accompany lower loss and perplexity in Section 3.2 demonstrate that CASTLE not only lowers perplexity but also translates these gains into stronger performance on diverse downstream tasks.
Table 2: Evaluation results (0-shot) for downstream tasks of different model scales. Each model is pretrained on FineWeb-Edu for 50B tokens. All values denote accuracy in percentage (%). The higher accuracy values are shown in bold. Hella.=HellaSwag, Wino.=Winograde.
Model Name ARC-C ARC-E BoolQ Hella. MMLU OBQA PIQA Wino. Avg.
Baseline-S $\bm{26.71}$ 54.76 52.51 35.78 22.89 30.40 63.98 $\bm{52.57}$ 42.45
CASTLE-S 26.19 $\bm{56.69}$ $\bm{59.85}$ $\bm{36.28}$ $\bm{23.00}$ $\bm{31.60}$ $\bm{64.25}$ 52.25 $\bm{43.76}$
Baseline-M 28.58 60.90 53.61 43.01 23.21 33.40 $\bm{67.95}$ 50.91 45.20
CASTLE-M $\bm{30.20}$ $\bm{61.36}$ $\bm{58.01}$ $\bm{43.24}$ $\bm{25.34}$ $\bm{34.60}$ $\bm{67.95}$ $\bm{52.64}$ $\bm{46.67}$
Baseline-L $\bm{32.59}$ 65.07 57.49 47.45 23.57 32.60 70.51 50.75 47.50
CASTLE-L 32.34 $\bm{65.15}$ $\bm{57.65}$ $\bm{47.87}$ $\bm{24.51}$ $\bm{35.60}$ $\bm{70.78}$ $\bm{53.51}$ $\bm{48.43}$
Baseline-XL 33.79 66.62 61.04 51.40 $\bm{26.72}$ 36.20 $\bm{72.58}$ 54.06 50.30
CASTLE-XL $\bm{35.32}$ $\bm{67.51}$ $\bm{62.81}$ $\bm{52.15}$ 23.74 $\bm{37.00}$ 70.67 $\bm{56.59}$ $\bm{50.72}$
Table 3: Evaluation results (5-shot) for downstream tasks of different model scales. The higher accuracy values are shown in bold. All values denote accuracy in percentage (%). Each model is pretrained on FineWeb-Edu for 50B tokens. Hella.=HellaSwag, Wino.=Winograde.
Model Name ARC-C ARC-E BoolQ Hella. MMLU OBQA PIQA Wino. Avg.
Baseline-S 25.68 $\bm{54.97}$ 56.09 33.81 $\bm{25.54}$ 28.20 63.98 $\bm{52.57}$ 42.60
CASTLE-S $\bm{26.02}$ 54.25 $\bm{57.13}$ $\bm{35.24}$ 25.22 $\bm{29.80}$ $\bm{64.53}$ 50.99 $\bm{42.90}$
Baseline-M 31.06 62.46 48.47 42.83 $\bm{25.22}$ 33.00 68.39 51.78 45.40
CASTLE-M $\bm{32.17}$ $\bm{64.06}$ $\bm{54.62}$ $\bm{43.47}$ $\bm{25.22}$ $\bm{33.80}$ $\bm{69.48}$ $\bm{52.49}$ $\bm{46.91}$
Baseline-L 33.36 63.64 $\bm{59.24}$ 46.16 $\bm{26.82}$ 33.40 69.53 $\bm{54.06}$ 48.28
CASTLE-L $\bm{37.37}$ $\bm{67.89}$ 50.95 $\bm{47.71}$ 26.11 $\bm{34.20}$ $\bm{70.18}$ $\bm{54.06}$ $\bm{48.56}$
Baseline-XL 35.58 65.78 61.07 50.84 $\bm{26.71}$ 36.20 $\bm{71.27}$ 52.72 50.02
CASTLE-XL $\bm{39.08}$ $\bm{70.24}$ $\bm{62.60}$ $\bm{51.63}$ 24.16 $\bm{37.40}$ 71.00 $\bm{58.33}$ $\bm{51.80}$
4 Conclusion
We introduced CAuSal aTtention with Lookahead kEys (CASTLE), a novel attention mechanism that continually updates keys as the context evolves. This design allows each key representation to incorporate more recent information at every prediction step while strictly preserving the autoregressive property. Although CASTLE is defined recurrently, we derived a mathematical equivalence that eliminates the need to explicitly materialize lookahead keys at each position, enabling efficient parallel training. Experimental results on language modeling demonstrate that CASTLE consistently outperforms standard causal attention, achieving lower perplexity and stronger downstream performance.